OpenStudy (anonymous):

line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2). Choose one answer. a. y = two thirdsx − 10 b. y = two thirdsx + 10 c. y = −three halvesx − 10 d. y = −three halvesx + 10

OpenStudy (gebooors):

First Solve y from equation 2x -3y = 3 Them, if another line is perpendicular with this one, K1 * k2 = -1

OpenStudy (anonymous):

3y = 2x - 3 y = (2/3)x -1 so slope is k1 = (2/3) perpendicular will be y = (-3/2)x + b find b by evaluating this new line for point (-8,2)